OpenStudy (anonymous):

t^2+6t-7. Is their a 2 numbers that equael product -7 and sum 6. Is it just me or thats not possible?

OpenStudy (kinggeorge):

That's just you :) There are indeed two numbers that multiply to be -7, and add to be 6. Think about it a bit and I'm sure you'll get it.

OpenStudy (anonymous):

Well the equation is set up as (a+1) (a-7) thats what I got. but 1-7=-6 so that doesnt work.

OpenStudy (anonymous):

remember that - and + in front of the numbers

OpenStudy (anonymous):

Wrong way around mate

OpenStudy (kinggeorge):

That's very close.

OpenStudy (anonymous):

ok got it switched them. Now it works.

OpenStudy (kinggeorge):

Excellent!

OpenStudy (anonymous):

Thanks lol I cant believe i cudnt figure tthat out

OpenStudy (kinggeorge):

No problem. We've all had something obvious pass us by before :P

OpenStudy (anonymous):

would the solution set be written as (a+7) (a-1) or just {7,-1}

OpenStudy (anonymous):

are you factorizing? normally it wuold be written as full.

OpenStudy (unklerhaukus):

you could always use the quadratic formula \[t=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

(a+7)(a-1) that is.... you will need to use the full factors to use in equations.

OpenStudy (kinggeorge):

The solution set would be {-7,1}. Note that you have -7 and +1 because if \(a+7=0\), then \(a=-7\). Similarly, \(a-1=0\implies a=1\).

OpenStudy (anonymous):

@Ben103 sorry the questions look like what I'm doing at school at the moment which is factorizing :( I could be wrong

OpenStudy (anonymous):

@KingGeorge knows more than me i'm sure! lol I'm only doing high school maths

OpenStudy (kinggeorge):

As another tip, if you're asked for a solution SET, your final answer should look like {_,_,_,...,_}. I.e., you have a list of things in the curly brackets.

OpenStudy (anonymous):

Ok thanks I understand it now. :)

OpenStudy (kinggeorge):

You're welcome.